OpenStudy (anonymous):

What is the rate of change from x=0 to x=pi/2?

OpenStudy (anonymous):

That link is not working for me

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

What is the rate of change from x = 0 to x = pi over 2? trig graph with points at 0, negative 4 and pi over 2, 0 and pi, 4 and 3 pi over 2, 0 and 2 pi, negative 4

jimthompson5910 (jim_thompson5910):

`trig graph with points at 0, negative 4 ... pi over 2, 0 ` so we know (0,-4) and (pi/2, 0) are points on this graph

OpenStudy (danjs):

the average rate is the slope of a line connecting those two endpoints

jimthompson5910 (jim_thompson5910):

you simply need to find the slope of the line through the two points (0,-4) and (pi/2, 0)

OpenStudy (anonymous):

I can find the slope fairly well, it's just with the pi in the equation, Idk what to do:(

OpenStudy (anonymous):

y1-y2/x1-x2=-4-0/0-pi right?

jimthompson5910 (jim_thompson5910):

I think you mean pi/2 instead of just pi? right?

OpenStudy (anonymous):

oh yes! Sorry

jimthompson5910 (jim_thompson5910):

so up top we have y2 - y1 = 0 - (-4) = 0 + 4 = 4 in the bottom we have x2 - x1 = pi/2 - 0 = pi/2

jimthompson5910 (jim_thompson5910):

now you just need to simplify \[\Large \frac{4}{\pi/2}\]

OpenStudy (anonymous):

8/pi?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Yay! Can I ask another/

OpenStudy (anonymous):

Compare the functions below: f(x) = 2x2 + 3x − 4 g(x) trig graph with points at 0, 0 and pi over 2, 3 and pi, 0 and 3 pi over 2, negative 3 h(x) x y −3 −8 −2 −6 −1 −4 0 −2 1 0 2 2 3 4 Which function has the largest y-intercept?

OpenStudy (anonymous):

would the answer be g(x)?

jimthompson5910 (jim_thompson5910):

I'm guessing g(x) is the graph being described with the set of points?

OpenStudy (anonymous):

yes:)

jimthompson5910 (jim_thompson5910):

what is the y intercept of f(x) ? g(x) ? h(x) ?

OpenStudy (anonymous):

when x=0 right?

jimthompson5910 (jim_thompson5910):

correct, what is the y value when x = 0 ?

OpenStudy (anonymous):

f(x)=2(0)^2+3(0)-4=-4 g(x)=0 h(x)-2

jimthompson5910 (jim_thompson5910):

all three look good. I'm assuming you meant to say `h(x) = -2 when x = 0` for that last part

OpenStudy (anonymous):

yes hahah :)

OpenStudy (anonymous):

So the largest y-intercept would be g(x) rigjt?

jimthompson5910 (jim_thompson5910):

yes since 0 is the largest of the three outputs (-4,0,-2)

OpenStudy (anonymous):

Thanks so much! You're awesome!

OpenStudy (anonymous):

I have more questions, so I'll open a new thread. Please come help again :D